OpenStudy (anonymous):

Need help, please explain, don't just give the answer ^_^ 7y – x2 + 3x −17x – 5y + 2x2 A. −13x2 + 1y B. x2 – 14x + 2y C. 2x2 – 12xy D. 3x2 + 20x + 12y

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@zepdrix

pooja195 (pooja195):

\[\huge~\rm~\bf~7y – x^2 + 3x −17x – 5y + 2x^2 \] combine like terms

pooja195 (pooja195):

\(\huge\rm\color{red}{(−x^2+2x^2)+(3x−17x)+(7y−5y)}\)

pooja195 (pooja195):

-1+2=?

OpenStudy (anonymous):

Sorry Openstudy wasn't working. I have a question, does it matter which order you put them as?

pooja195 (pooja195):

Yes 

pooja195 (pooja195):

^example

pooja195 (pooja195):

The like terms must be together so that they can be combined to simplfy the expression

jimthompson5910 (jim_thompson5910):

Imagine you're a doctor working with an x-ray. You have a patient who needs 2 x-rays. You have another patient that needs 3 x-rays. In total, 2+3 = 5 x-rays are done. If you wanted to make a shorthand note of the x-rays, you could use algebraic notation For the first patient who needs 2 x-rays, you could write 2x For the first patient who needs 3 x-rays, you could write 3x Adding 2x to 3x gives you 2x+3x = 5x This example shows how combining like terms works. Now imagine we throw in some other variable y. Say y for yogurt. If we have 3x+10y we CANNOT add them because we can't say 3x+10y = 13x or 13y or 13xy. It makes no sense to say any of those three things. The two things are completely separate and can't be combined. So we just leave 3x+10y as it is Anyways, that's just an example that hopefully helps. I'll let @pooja195 continue

OpenStudy (anonymous):

Thanks Jim for explaining

OpenStudy (anonymous):

But what I was asking was does it matter which order you put the group of like terms? For example, instead of what you put, what if you did it in a different order? (with the like terms still together)

pooja195 (pooja195):

It does matter. You can put them in a different order but you need to move the sign with it. Like -x^2+2x^2 can also be written as 2x^2-x^2 and still has the same meaning \[\huge~\rm~\bf~7y – x^2 + 3x −17x – 5y + 2x^2\]

OpenStudy (anonymous):

Still not sure how to put them in order tho

pooja195 (pooja195):



pooja195 (pooja195):

Hint: Look at the signs

jimthompson5910 (jim_thompson5910):

it doesn't matter because x+y = y+x in other words, you can add terms in any order

OpenStudy (anonymous):

Wait then why can't it be 7y-5y?

pooja195 (pooja195):

You can write it like that

OpenStudy (anonymous):

Meaning 7y-5y first

pooja195 (pooja195):

That's fine the order doesn't matter unless you have negatives then u just need to becareful to move the negative sign

pooja195 (pooja195):

7y-5y is fine

OpenStudy (anonymous):

Ok so you can continue with the question

pooja195 (pooja195):

From here you just need to add/subtract.. \(\huge\rm\color{red}{(−x^2+2x^2)+(3x−17x)+(7y−5y)}\) -1+2=?

OpenStudy (anonymous):

1?

pooja195 (pooja195):

Great so now we have From here you just need to add/subtract.. \(\huge\rm\color{red}{(x^2)+(3x−17x)+(7y−5y)}\) 3x-17x=?

OpenStudy (anonymous):

-14?

pooja195 (pooja195):

Perfect! :) \(\huge\rm\color{red}{(x^2)+(-14x)+(7y−5y)}\) 7y-5y=?

OpenStudy (anonymous):

2y

pooja195 (pooja195):

Perfect! :) \(\huge\rm\color{red}{(x^2)+(-14x)+(2y)}\) Now remove the parentheses + sign goes away since it is -14 \(\huge\rm\color{red}{x^2-14x+2y}\)

OpenStudy (anonymous):

Ok, thanks Pooja! (And Jim)

pooja195 (pooja195):

yw ^.^